OpenStudy (yorel_1999):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(4)] 2

OpenStudy (xguardians):

Can you please fix your equations? Is that f(x) = x^2 + 1? What is the Find [f(4)]^2?

OpenStudy (yorel_1999):

Im Sorry. That's Exactly How The Equation Was Written. And I'm Really Bad In This Subject. :/

OpenStudy (shamim):

given \[f(x)=x ^{2}+1\]

OpenStudy (shamim):

so \[f(4)=4^{2}+1\]

OpenStudy (shamim):

right?

OpenStudy (yorel_1999):

I believe so.

OpenStudy (shamim):

\[4^{2}=?\]

OpenStudy (yorel_1999):

If i give you the answer choices that are attached to this question, will it help?

OpenStudy (shamim):

later

OpenStudy (yorel_1999):

Okay

OpenStudy (shamim):

plz try to tell \[4^{2}=?\]

OpenStudy (shamim):

just try, i wanna help u if u r unable to say

OpenStudy (shamim):

can u tell me \[1^{2}=?\]

OpenStudy (yorel_1999):

I really need help. I dont know what to do. I was given This question and I don't know how to solve it. I'm sorry that I can't really tell you.

OpenStudy (shamim):

ok

OpenStudy (shamim):

u know \[1^{2}=1 \times 1=?\]

OpenStudy (yorel_1999):

1x1 is 1^2

OpenStudy (shamim):

can u tell me \[1+2=?\]

OpenStudy (yorel_1999):

3...

OpenStudy (shamim):

good!!

OpenStudy (shamim):

can u tell me \[1 \times 1=?\]

OpenStudy (yorel_1999):

1...

OpenStudy (shamim):

right u r !!

OpenStudy (shamim):

the rule of math is \[1^{2}=1 \times 1\]

OpenStudy (shamim):

now tell me plz \[1^{2}=?\]

OpenStudy (shamim):

dont feel hegitation

OpenStudy (shamim):

response plz

OpenStudy (shamim):

tell me \[1^{2}=?\]

OpenStudy (yorel_1999):

1x1